Kids Web

A World Wide Web Digital Library for Schoolkids

THE WORLD WIDE WEB

The ambitious goal of the World Wide Web project is to make available all of human knowledge through global computer networks acting as a vast distributed electronic library. The Web allows multimedia information (text, audio, images, video) to be be accessed over the Internet with a simple point-and-click interface by using a Web browser such as Mosaic or Netscape.

The Web contains enormous amounts of multimedia information, and is growing incredibly rapidly. However only a small fraction of that information is useful, intelligible and interesting to schoolchildren. Finding this information using standard Web digital libraries and search methods can be very time-consuming and frustrating, especially for schoolkids. Ideally the students should be presented with a subset of the Web that is very simple to navigate, and contains information targeted at the K-12 level. This is the goal of Kids Web. Each of the subject sections contains a list of links to information that is understandable and interesting to schoolkids.

THE HISTORY AND FUTURE OF KIDS WEB

Kids Web grew out of a computer multimedia class based on the World Wide Web and Mosaic which was taught to 8th graders as part of the 1994 Young Scholars Program at Syracuse University. It is being developed for The Living Textbook project, which applies high-performance computing and communications to K-12 education, and is funded by New York State. By early 1995, a number of schools in the state will be connected to the Northeast Parallel Architectures Center (NPAC) at Syracuse University by NYNet, a state-of-the-art fast optical fiber network.

Currently the Information Superhighway more closely resembles a dirt road, so accessing information on the Web can be very slow, especially material from computers in other parts of the world. This is a major problem, especially for children with low attention spans. We will use Web caching technology to pre-fetch the documents accessible via Kids Web and store them on a centralized server at NPAC. This will allow the information to be accessed much more rapidly, especially by students connected to NYNet. It will also enable schools to easily address the issue of eliminating access to undesirable material on the Internet, by allowing access only to the cached Kids Web resources.

In order for Kids Web to fulfill the goal of being a comprehensive digital library for kids, it must be managed in a distributed fashion, so that teachers (and students) from all over the world can suggest links to relevant Web documents, and each subject listing will be edited by educators who specialize in that field. If you are interested in being an editor for a subject section, or want to suggest a document to be linked to Kids Web, please email kidsweb@npac.syr.edu. Comments and inquiries are also welcome.

WHERE TO FIND KIDS WEB

Kids Web is now available by directing your World Wide Web browser (e.g. Mosaic or Netscape) to http://www.npac.syr.edu/textbook/kidsweb/

Paul Coddington

Northeast Parallel Architectures Center Syracuse University kidsweb@npac.syr.edu